If $A$ and $B$ are independent events and $P(A) = 0.4$,$P(A \cup B) = 0.6$,and $P(B) = p$,then the value of $p$ is . . . . . . .

  • A
    $1/2$
  • B
    $3/4$
  • C
    $1/3$
  • D
    $5/6$

Explore More

Similar Questions

An urn contains $4$ red and $5$ white balls. Two balls are drawn one after the other without replacement,then the probability that both the balls are red is

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

Bag $A$ contains $4$ white and $2$ black balls, bag $B$ contains $3$ white and $3$ black balls, and bag $C$ contains $2$ white and $4$ black balls. If a bag is chosen at random and a ball is chosen at random from it, then the probability that the ball drawn is black is

$A$ pandemic has been spreading all over the world. The probabilities are $0.7$ that there will be a lockdown,$0.8$ that the pandemic is controlled in one month if there is a lockdown,and $0.3$ that it is controlled in one month if there is no lockdown. The probability that the pandemic will be controlled in one month is

Given two independent events $A$ and $B$ such that $P(A) = 0.3$ and $P(B) = 0.6$,find $P(A \text{ and } B)$.

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo